# Preserving curvature lower bounds when Ricci flowing non-smooth initial data

### v1

@@ 1. Introduction
    $M=M^{3},$ ${\rm R}(g(0))>0$ and ${\rm Ric}(g(0))\geq\varepsilon{\rm R}(g(0))$implies ${\rm R}(g(t))>0$ and ${\rm Ric}(g(t))\geq\varepsilon{\rm R}(g(t))$ for all $t\in[0,T]$ : Hamilton [Ham82] (Lott [Lot19]).
@@
Such spaces may have singularities, either of the metric type or in the topological sense. A metric space $(M,d)$ is smooth at a point $p\in M$ if there is a neighbourhood $U$ of $p$ in $M$ and a homeomorphism $\varphi:U\subseteq M\to\mathbb{B}_{r}(0)\subseteq\mathbb{R}^{n},$ and a smooth Riemannian metric $g$ on $\mathbb{B}_{r}(0)$ such that the push forward $\tilde{d}$ of $d$, $\tilde{d}(x,y):=d(\varphi^{-1}(x),\varphi^{-1}(y))$ agrees with the metric $\hat{d}$ induced by $(\mathbb{B}_{r}(0),g)$ on $\mathbb{B}_{s}(0)$ for some $s>0$.
@@ 2. Ricci flows of non-smooth data
iii) We use the following notation from [BGP92]. Let $(X_{i},d_{i}),(X,d_{X})$ be metric spaces, $i\in\mathbb{N},$ and $x_{i}\in X_{i},$ $x\in X$ for all $i\in\mathbb{N}$. We say $(X_{i},d_{i},x_{i})$ converges in the pointed Gromov-Hausdorff sense to $(X,d_{X},x)$ if there exists $\varepsilon(i),R(i)\in\mathbb{R}^{+}$ with $\varepsilon(i)\to 0$ and $R(i)\to\infty$ as $i\to\infty$ and maps $F_{i}:B_{d_{i}}(x_{i},R(i))\to B_{d_{X}}(x,R(i))$ with $F_{i}(x_{i})=x$ such that
$|d_{X}(F_{i}(y),F_{i}(z))-d_{i}(y,z)|\leq\varepsilon(i)$
for all $i\in\mathbb{N}$ for all $x,y\in B_{d_{i}}(x_{i},R(i)),$ and $B_{d_{X}}(x,R(i)-\varepsilon(i))\subseteq F_{i}(B_{d_{i}}(x_{i},R(i)))$.
@@
    $g(0)\in C^{0}(M),$ $(M,h)$ smooth, complete , connected, $\sup_{M}|{\rm Rm}(h)|<\infty$ , and assume $h$ is $(1+\varepsilon(n))$-fair to $g(0)$ for a small $\varepsilon(n)>0$. Then there is a Ricci flow solution coming out of $(M,d(g(0))).$ Simon [Sim01] and [Sim02] (see also Koch/Lamm [KL12, KL15]). See Section 3.
@@
    $\int_{M}(|{{}^{h}\nabla}g(0)|^{2}+|{{}^{h}\nabla}^{2}g(0)|^{2})d{\rm vol}_{h}<\infty$ (2.1)
@@
    $(X,d_{X},x)$ is the pointed Gromov-Hausdorff limit of a sequence of smooth, three dimensional Riemannian manifolds $(M^{3}_{i},g_{i}(0),x_{i})$ which are uniformly non-collapsed, that is ${\rm vol}(B_{g_{i(0)}}(x,1))\geq v_{0}>0$ for some $v_{0}>0$ for all $x\in M_{i}$, and the metrics satisfy ${\rm Ric}(g_{i}(0))+kg_{i}(0)\geq 0$ for some $k\in\mathbb{R}$ or $\sec(g_{i}(0))+k\geq 0$ for some $k\in\mathbb{R}$. Then there is a smooth Ricci flow $(M,g(t))_{t\in(0,T)}$ coming out of $(X,d_{X}).$ Simon [Sim12] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|<\infty$ for all $i\in\mathbb{N}$ , Simon/Topping [ST22, ST21], Hochard [Hoc16] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|=\infty$ for all or some $i\in\mathbb{N}.$
@@
    $(X,d_{X},x)$ is the pointed Gromov-Hausdorff limit of a sequence of smooth, connected, complete Riemannian manifolds $(M^{n}_{i},g_{i}(0),x_{i}))$ which are uniformly non-collapsed, that is ${\rm vol}(B_{g_{i}}(x,1))\geq v_{0}>0$ for some $v_{0}>0$ for all $x\in M_{i}$, and the metrics satisfy $\mathcal{R}(g_{i}(0))+kI(g_{i}(0))\in\mathcal{C}$ for some $k\in\mathbb{R}$ where $\mathcal{C}$ is one of the curvature conditions $\mathcal{C}_{CO},\mathcal{C}_{2C0},\mathcal{C}_{IC1},\mathcal{C}_{IC2},$ in the list of curvature conditions 2.3 given below. Then there is a smooth Ricci flow $(M,g(t))_{t\in(0,T)}$ coming out of $(X,d_{X}):$ Bamler/Cabezas-Rivas/Wilking, [BCRW19] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|<\infty$ for all $i\in\mathbb{N},$ and Bamler/Cabezas-Rivas/Wilking, [BCRW19] Hochard [Hoc19], Lai [Lai19], in the case $\sup_{M}|{\rm Rm}(g_{i}(0))|=\infty$ for some or all $i\in\mathbb{N}$ : See Section 4 for more details/references.
@@
    $(X,d_{X},x)$ is the pointed Gromov-Hausdorff limit of a sequence of smooth, connected complete Kähler manifolds $(M^{n}_{i},g_{i}(0),x_{i})$ which are uniformly non-collapsed, that is ${\rm vol}(B_{g_{i}(0)}(x,1))\geq v_{0}>0$ for some $v_{0}>0$ for all $x\in M_{i}$, and the metrics satisfy $\mathcal{R}(g_{i}(0))+kI(g_{i}(0))\in\mathcal{C}$ for some $k\in\mathbb{R}$ where $\mathcal{C}$ is the condition $\mathcal{C}_{HB}$ of 2.3. Then there is a smooth Kähler Ricci flow $(M,g(t))_{t\in(0,T)}$ coming out of $(X,d_{X}).$ Bamler/Cabezas-Rivas/Wilking [BCRW19], in the case $\sup_{M}|{\rm Rm}(g_{i}(0))|<\infty$ for all $i\in\mathbb{N},$ Lee/Tam [LT21b] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|=\infty$ for some or all $i\in\mathbb{N}.$
@@
    Note that if $M$ is compact in (iii) or (iv), $g\in W^{k,l}_{glob}(M^{n},h)$ if and only if $g\in W^{k,l}_{loc}(M^{n})$ where $g\in W^{k,l}_{loc}(M^{n})$ if and only we can find a covering of $M$ by coordinate charts $(U_{\alpha})_{\alpha=1}^{N},$ $\varphi_{\alpha}:U_{\alpha}\to V_{\alpha}\subseteq\mathbb{R}^{n}$ with inverse $\psi_{\alpha}$ for which, $(g_{\alpha})_{ij}(x):=g(\psi_{\alpha}(x))(D\psi_{\alpha}(x)(e_{i}(x)),D\psi_{\alpha}(x)(e_{j}(x)))$ for $x\in V_{\alpha}$ satisfies $(g_{\alpha})_{ij}\in W^{k,l}(V_{\alpha})$ for each $\alpha\in\{1,\ldots,N\}$ for each $i,j\in\{1,\ldots,n\},$ $(k,l)=(1,n)$ in the case (iv) above, or $(k,l)=(2,2),$ in the case (iii) above, and $W^{k,l}(V_{\alpha})$ refers to the usual definition of a Sobolev space.
@@
Note: The conditions $S\in\mathcal{C}^{+}_{IC},$ respectively $S\in\mathcal{C}^{+}_{IC1},$ respectively $S\in\mathcal{C}^{+}_{IC2},$ are generally known in the literature as "$S$ has positive isotropic curvature ", or $S$ is PIC, respectively "$S$ is PIC1", respectively "$S$ is PIC2". The condition PIC was introduced in the paper [MM88], the conditions PIC1, PIC2 in the paper [BS09].
@@ 3. $C^{0}$ Limits of smooth Riemannian manifolds with curvature lower bounds
In the papers [Sim01] and [Sim02] (in [Sim02] some corrrections to [Sim01] were made) it was shown for any $C^{0}$ Riemannian metric $g(0)$ on a closed manifold $M$, and any smooth metric $h$, which satisfies $\frac{1}{(1+\varepsilon(n))}h\leq g(0)\leq(1+\varepsilon(n))h$ for a small constant $\varepsilon(n)>0$ (such an $h$ may be obtained by mollifying the metric $g(0)$), then there exists a solution $g(t)_{t\in(0,T)}$ to the so called Ricci DeTurck flow with background metric $h$, such that $|g(t)-g(0)|_{h}\to 0$ as $t\searrow 0,$ where $T=T(n,\sup_{M}|{\rm Rm}(h)|_{h})>0.$ The Ricci DeTurck flow was first studied by DeTurck in the paper [DeT83]. For simplicity we call Ricci DeTurck flow with background metric $h$ the $h$-Ricci DeTurck flow. For an interval $I$, $(M,g(t))_{t\in I}$ is a solution to $h$-Ricci DeTurck flow if it solves the equation
@@
( [Sim01] and [Sim02] ) For any $n\in\mathbb{N}$ there exists an $\varepsilon(n)>0$ such that the following holds. Let $(M^{n},h)$ be a smooth, connected, complete $n$-dimensional Riemannian manifold such that $K_{j}:=\sup_{M}|\nabla^{j}{\rm Rm}(h)|<\infty$ for all $j\in\mathbb{N}_{0}.$ Assume further that $g(0)$ is a $C^{0}$ Riemannian metric on $M,$ for which $\frac{1}{(1+\hat{\varepsilon})}h\leq g(0)\leq(1+\hat{\varepsilon})h$ holds, where $0<\hat{\varepsilon}\leq\varepsilon(n).$ Then there exists a $T=T(n,\hat{\varepsilon},K_{0})>0$, $T$ being monotone increasing in the second variable and decreasing in the third variable, and a smooth complete solution $(M,g(t))_{t\in(0,T)}$ to (3.1) with initial value $g(0)$, in the sense that $\sup_{K}|g(t)-g(0)|_{g(0)}\to 0$ for any compact $K\subseteq M,$ as $t\searrow 0,$ and the following estimates hold:
$\frac{1}{(1+2\hat{\varepsilon})}h\leq g(t)\leq(1+2\hat{\varepsilon})h$
@@
If $(M,\hat{h})$ is a smooth, complete, connected manifold with $\hat{K}_{0}:=\sup_{M}|{\rm Rm}(\hat{h})|<\infty,$ and $g(0)$ a $C^{0}$ Riemannian metric such that $g(0)$ is $1+\frac{\varepsilon(n)}{2}$ fair to $\hat{h}$, then we can find, using Theorem 1.2 of Shi ([Shi89]), an $h$ satisfying the conditions of the Theorem.
@@
If $M$ is closed, smooth, and $g(0)$ a given $C^{0}$ Riemannian metric, then an $h$ and $\hat{\varepsilon}$ as in the statement of the theorem always exist.
@@
(Lamm/Simon [LS23]) There exists an $\varepsilon_{1}(n)>0$ with the following properties. Let $M^{n}$ be smooth, $g(0),h,$ $\hat{\varepsilon},\varepsilon(n)$ be as in in the statement of Theorem 3.1 and assume further that $inj(M,h)\geq i_{0}>0,$ and $\varepsilon(n)<\varepsilon_{1}(n)$. Then there exists a smooth, complete, Ricci flow solution $\ell(t)_{t\in(0,T)}$ with $\sup_{K}|d(\ell(t))-d_{0}|\to 0$ for any compact $K\subseteq M,$ as $t\searrow 0,$ where $d_{0}$ is a metric isometric to $d(g(0))$. The solution $\ell(t)$ is a Ricci flow related solution to the Ricci DeTurck solution $g(t)$ from Theorem 3.1, and can be written as $\ell(t):=(\Phi_{t})^{*}g(t),$ where $\Phi:M\times(0,T)\to M$ is a solution to (3.4) with $S=\frac{T}{2},$ and $\Phi_{t}:=\Phi(\cdot,t)$ are diffeomorphisms for $t\in(0,T),$ $\Phi_{t}\to\Phi_{0}$ locally uniformly as $t\searrow 0$ where $\Phi_{0}$ is a homeomorphism. The metric $d_{0}$ is given by $d_{0}:=\Phi_{0}^{*}(d(g(0)))$
@@
We let $\Phi:M\times(0,T)\to M$ be the solution to (5.8) with $S=\frac{T}{2}$. The paper of Lamm/Simon then shows (see Theorem 2.6 in [LS23]) that $\Phi(\cdot,t)\to\Phi_{0}$ locally uniformly as $t\searrow 0$ where $\Phi_{0}$ is a homeomorphism. Now using that $d(g(t))\to d_{g(0)}$ locally uniformly, we see that $d_{\ell(t)}(x,y)=d(g(t))(\Phi_{t}(x),\Phi_{t}(y))$ satisfies $d_{\ell(t)}\to(\Phi_{0})^{*}d_{g(0)}=:d_{0}$ locally uniformly. The completeness may be seen as followed: For fixed $t$ let $(x_{i})_{i\in\mathbb{N}}$ be a Cauchy sequence with respect to $d(\ell(t))$. We have, $d_{\ell(t)}(x_{i},x_{j})\leq\varepsilon(N)$ for all $i,j\geq N$ where $\varepsilon(N)\to 0$ as $N\to\infty.$ Hence $d_{g(t)}(\Phi_{t}(x_{i}),\Phi_{t}(x_{j}))\leq\varepsilon(N)$ for all $i,j\geq N$. Since $g(t)$ is $a$-fair to $h$, we know that $g(t)$ is complete. Hence, $\Phi_{t}(x_{i})$ converges with respect to $g(t)$ to an $p$ as $i\to\infty$ which we can write uniquely as $p=\Phi_{t}(q),$ since $\Phi_{t}$ is a smooth diffeomorphism. Hence $(x_{i})_{i\in\mathbb{N}}$ converges, as $i\to\infty$, with respect to $\ell(t)$ to $q$. ∎
@@
Let $(M,g(0))$ be $C^{0}$ and $h$ be fixed, and satisfy $\sup_{M}|\nabla^{j}{\rm Rm}(h)|_{h}=:K_{j}<\infty.$ Let $\hat{\varepsilon}\leq\varepsilon(n)$ satisfy $\frac{1}{(1+\hat{\varepsilon})}h\leq g(0)\leq(1+\hat{\varepsilon})h$ and let $g(t)_{t\in(0,T)}$ be the solution to (3.1) from Theorem 3.1. Then there exists $\hat{T}(n,h,\hat{\varepsilon})>0,\hat{T}\leq T$ such that
$|{\rm Rm}(g(t))|\leq\frac{\psi(n,\hat{\varepsilon})}{t}$
for all $t\in(0,\hat{T}(n,h,\hat{\varepsilon})),$ where $\psi(n,\hat{\varepsilon})\to 0$ as $\hat{\varepsilon}\to 0$. Hence the same is true for any Ricci flow related solution. In particular, if $inj(M,h)>0,$ then the Ricci flow related solution obtained in Theorem 3.4 satisfies this estimate also:
$|{\rm Rm}(\ell(t))|\leq\frac{\psi(n,\hat{\varepsilon})}{t}$
for all $t\in(0,\hat{T}(n,h,\hat{\varepsilon})),$ where $\psi(n,\hat{\varepsilon})\to 0$ as $\hat{\varepsilon}\to 0$.

### v2

@@ 1. Introduction
    $M=M^{3},$ ${\rm R}(g(0))>0$ and ${\rm Ric}(g(0))\geq\epsilon{\rm R}(g(0))$implies ${\rm R}(g(t))>0$ and ${\rm Ric}(g(t))\geq\epsilon{\rm R}(g(t))$ for all $t\in[0,T]$ : Hamilton [Ham82] (Lott [Lot19]).
@@
Such spaces may have singularities, either of the metric type or in the topological sense. A metric space $(M,d)$ is smooth at a point $p\in M$ if there is a neighbourhood $U$ of $p$ in $M$ and a homeomorphism $\phi:U\subseteq M\to\mathbb{B}_{r}(0)\subseteq\mathbb{R}^{n},$ and a smooth Riemannian metric $g$ on $\mathbb{B}_{r}(0)$ such that the push forward $\tilde{d}$ of $d$, $\tilde{d}(x,y):=d(\phi^{-1}(x),\phi^{-1}(y))$ agrees with the metric $\hat{d}$ induced by $(\mathbb{B}_{r}(0),g)$ on $\mathbb{B}_{s}(0)$ for some $s>0$.
@@ 2. Ricci flows of non-smooth data
iii) We use the following notation from [BGP92]. Let $(X_{i},d_{i}),(X,d_{X})$ be metric spaces, $i\in\mathbb{N},$ and $x_{i}\in X_{i},$ $x\in X$ for all $i\in\mathbb{N}$. We say $(X_{i},d_{i},x_{i})$ converges in the pointed Gromov-Hausdorff sense to $(X,d_{X},x)$ if there exists $\epsilon(i),R(i)\in\mathbb{R}^{+}$ with $\epsilon(i)\to 0$ and $R(i)\to\infty$ as $i\to\infty$ and maps $F_{i}:B_{d_{i}}(x_{i},R(i))\to B_{d_{X}}(x,R(i))$ with $F_{i}(x_{i})=x$ such that
$|d_{X}(F_{i}(y),F_{i}(z))-d_{i}(y,z)|\leq\epsilon(i)$
for all $i\in\mathbb{N}$ for all $x,y\in B_{d_{i}}(x_{i},R(i)),$ and $B_{d_{X}}(x,R(i)-\epsilon(i))\subseteq F_{i}(B_{d_{i}}(x_{i},R(i)))$.
@@
    $g(0)\in C^{0}(M),$ $(M,h)$ smooth, complete , connected, $\sup_{M}|{\rm Rm}(h)|<\infty$ , and assume $h$ is $(1+\epsilon(n))$-fair to $g(0)$ for a small $\epsilon(n)>0$. Then there is a Ricci flow solution coming out of $(M,d(g(0))).$ Simon [Sim01] and [Sim02] (see also Koch/Lamm [KL12, KL15]). See Section 3.
@@
    $\int_{M}(|{{}^{h}\nabla}g(0)|^{2}+|{{}^{h}\nabla}^{2}g(0)|^{2})dV_{h}<\infty$ (2.1)
@@
    $(X,d_{X},x)$ is the pointed Gromov-Hausdorff limit of a sequence of smooth, three dimensional Riemannian manifolds $(M^{3}_{i},g_{i}(0),x_{i})$ which are uniformly non-collapsed, that is $V(B_{g_{i(0)}}(x,1))\geq v_{0}>0$ for some $v_{0}>0$ for all $x\in M_{i}$, and the metrics satisfy ${\rm Ric}(g_{i}(0))+kg_{i}(0)\geq 0$ for some $k\in\mathbb{R}$ or $\sec(g_{i}(0))+k\geq 0$ for some $k\in\mathbb{R}$. Then there is a smooth Ricci flow $(M,g(t))_{t\in(0,T)}$ coming out of $(X,d_{X}).$ Simon [Sim12] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|<\infty$ for all $i\in\mathbb{N}$ , Simon/Topping [ST22, ST21], Hochard [Hoc16] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|=\infty$ for all or some $i\in\mathbb{N}.$
@@
    $(X,d_{X},x)$ is the pointed Gromov-Hausdorff limit of a sequence of smooth, connected, complete Riemannian manifolds $(M^{n}_{i},g_{i}(0),x_{i}))$ which are uniformly non-collapsed, that is $V(B_{g_{i}}(x,1))\geq v_{0}>0$ for some $v_{0}>0$ for all $x\in M_{i}$, and the metrics satisfy $\mathcal{R}(g_{i}(0))+kI(g_{i}(0))\in\mathcal{C}$ for some $k\in\mathbb{R}$ where $\mathcal{C}$ is one of the curvature conditions $\mathcal{C}_{CO},\mathcal{C}_{2C0},\mathcal{C}_{IC1},\mathcal{C}_{IC2},$ in the list of curvature conditions 2.3 given below. Then there is a smooth Ricci flow $(M,g(t))_{t\in(0,T)}$ coming out of $(X,d_{X}):$ Bamler/Cabezas-Rivas/Wilking, [BCRW19] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|<\infty$ for all $i\in\mathbb{N},$ and Bamler/Cabezas-Rivas/Wilking, [BCRW19] Hochard [Hoc19], Lai [Lai19], in the case $\sup_{M}|{\rm Rm}(g_{i}(0))|=\infty$ for some or all $i\in\mathbb{N}$ : See Section 4 for more details/references.
@@
    $(X,d_{X},x)$ is the pointed Gromov-Hausdorff limit of a sequence of smooth, connected complete Kähler manifolds $(M^{n}_{i},g_{i}(0),x_{i})$ which are uniformly non-collapsed, that is $V(B_{g_{i}(0)}(x,1))\geq v_{0}>0$ for some $v_{0}>0$ for all $x\in M_{i}$, and the metrics satisfy $\mathcal{R}(g_{i}(0))+kI(g_{i}(0))\in\mathcal{C}$ for some $k\in\mathbb{R}$ where $\mathcal{C}$ is the condition $\mathcal{C}_{HB}$ of 2.3. Then there is a smooth Kähler Ricci flow $(M,g(t))_{t\in(0,T)}$ coming out of $(X,d_{X}).$ Bamler/Cabezas-Rivas/Wilking [BCRW19], in the case $\sup_{M}|{\rm Rm}(g_{i}(0))|<\infty$ for all $i\in\mathbb{N},$ Lee/Tam [LT21b] for the case $\sup_{M}|{\rm Rm}(g_{i}(0))|=\infty$ for some or all $i\in\mathbb{N}.$
@@
    Note that if $M$ is compact in (iii) or (iv), $g\in W^{k,l}_{glob}(M^{n},h)$ if and only if $g\in W^{k,l}_{loc}(M^{n})$ where $g\in W^{k,l}_{loc}(M^{n})$ if and only we can find a covering of $M$ by coordinate charts $(U_{\alpha})_{\alpha=1}^{N},$ $\phi_{\alpha}:U_{\alpha}\to V_{\alpha}\subseteq\mathbb{R}^{n}$ with inverse $\psi_{\alpha}$ for which, $(g_{\alpha})_{ij}(x):=g(\psi_{\alpha}(x))(D\psi_{\alpha}(x)(e_{i}(x)),D\psi_{\alpha}(x)(e_{j}(x)))$ for $x\in V_{\alpha}$ satisfies $(g_{\alpha})_{ij}\in W^{k,l}(V_{\alpha})$ for each $\alpha\in\{1,\ldots,N\}$ for each $i,j\in\{1,\ldots,n\},$ $(k,l)=(1,n)$ in the case (iv) above, or $(k,l)=(2,2),$ in the case (iii) above, and $W^{k,l}(V_{\alpha})$ refers to the usual definition of a Sobolev space.
@@
Note: The conditions $S\in\mathcal{C}^{+}_{IC},$ respectively $S\in\mathcal{C}^{+}_{IC1},$ respectively $S\in\mathcal{C}^{+}_{IC2},$ are generally known in the literature as ”$S$ has positive isotropic curvature ”, or $S$ is PIC, respectively ”$S$ is PIC1”, respectively ”$S$ is PIC2”. The condition PIC was introduced in the paper [MM88], the conditions PIC1, PIC2 in the paper [BS09].
@@ 3. $C^{0}$ Limits of smooth Riemannian manifolds with curvature lower bounds
In the papers [Sim01] and [Sim02] (in [Sim02] some corrrections to [Sim01] were made) it was shown for any $C^{0}$ Riemannian metric $g(0)$ on a closed manifold $M$, and any smooth metric $h$, which satisfies $\frac{1}{(1+\epsilon(n))}h\leq g(0)\leq(1+\epsilon(n))h$ for a small constant $\epsilon(n)>0$ (such an $h$ may be obtained by mollifying the metric $g(0)$), then there exists a solution $g(t)_{t\in(0,T)}$ to the so called Ricci DeTurck flow with background metric $h$, such that $|g(t)-g(0)|_{h}\to 0$ as $t\searrow 0,$ where $T=T(n,\sup_{M}|{\rm Rm}(h)|_{h})>0.$ The Ricci DeTurck flow was first studied by DeTurck in the paper [DeT83]. For simplicity we call Ricci DeTurck flow with background metric $h$ the $h$-Ricci DeTurck flow. For an interval $I$, $(M,g(t))_{t\in I}$ is a solution to $h$-Ricci DeTurck flow if it solves the equation
@@
( [Sim01] and [Sim02] ) For any $n\in\mathbb{N}$ there exists an $\epsilon(n)>0$ such that the following holds. Let $(M^{n},h)$ be a smooth, connected, complete $n$-dimensional Riemannian manifold such that $K_{j}:=\sup_{M}|\nabla^{j}{\rm Rm}(h)|<\infty$ for all $j\in\mathbb{N}_{0}.$ Assume further that $g(0)$ is a $C^{0}$ Riemannian metric on $M,$ for which $\frac{1}{(1+\hat{\epsilon})}h\leq g(0)\leq(1+\hat{\epsilon})h$ holds, where $0<\hat{\epsilon}\leq\epsilon(n).$ Then there exists a $T=T(n,\hat{\epsilon},K_{0})>0$, $T$ being monotone increasing in the second variable and decreasing in the third variable, and a smooth complete solution $(M,g(t))_{t\in(0,T)}$ to (3.1) with initial value $g(0)$, in the sense that $\sup_{K}|g(t)-g(0)|_{g(0)}\to 0$ for any compact $K\subseteq M,$ as $t\searrow 0,$ and the following estimates hold:
$\frac{1}{(1+2\hat{\epsilon})}h\leq g(t)\leq(1+2\hat{\epsilon})h$
@@
If $(M,\hat{h})$ is a smooth, complete, connected manifold with $\hat{K}_{0}:=\sup_{M}|{\rm Rm}(\hat{h})|<\infty,$ and $g(0)$ a $C^{0}$ Riemannian metric such that $g(0)$ is $1+\frac{\epsilon(n)}{2}$ fair to $\hat{h}$, then we can find, using Theorem 1.2 of Shi ([Shi89]), an $h$ satisfying the conditions of the Theorem.
@@
If $M$ is closed, smooth, and $g(0)$ a given $C^{0}$ Riemannian metric, then an $h$ and $\hat{\epsilon}$ as in the statement of the theorem always exist.
@@
(Lamm/Simon [LS23]) There exists an $\epsilon_{1}(n)>0$ with the following properties. Let $M^{n}$ be smooth, $g(0),h,$ $\hat{\epsilon},\epsilon(n)$ be as in in the statement of Theorem 3.1 and assume further that $inj(M,h)\geq i_{0}>0,$ and $\epsilon(n)<\epsilon_{1}(n)$. Then there exists a smooth, complete, Ricci flow solution $\ell(t)_{t\in(0,T)}$ with $\sup_{K}|d(\ell(t))-d_{0}|\to 0$ for any compact $K\subseteq M,$ as $t\searrow 0,$ where $d_{0}$ is a metric isometric to $d(g(0))$. The solution $\ell(t)$ is a Ricci flow related solution to the Ricci DeTurck solution $g(t)$ from Theorem 3.1, and can be written as $\ell(t):=(\Phi_{t})^{*}g(t),$ where $\Phi:M\times(0,T)\to M$ is a solution to (3.4) with $S=\frac{T}{2},$ and $\Phi_{t}:=\Phi(\cdot,t)$ are diffeomorphisms for $t\in(0,T),$ $\Phi_{t}\to\Phi_{0}$ locally uniformly as $t\searrow 0$ where $\Phi_{0}$ is a homeomorphism. The metric $d_{0}$ is given by $d_{0}:=\Phi_{0}^{*}(d(g(0)))$
@@
We let $\Phi:M\times(0,T)\to M$ be the solution to (5.8) with $S=\frac{T}{2}$. The paper of Lamm/Simon then shows (see Theorem 2.6 in [LS23]) that $\Phi(\cdot,t)\to\Phi_{0}$ locally uniformly as $t\searrow 0$ where $\Phi_{0}$ is a homeomorphism. Now using that $d(g(t))\to d_{g(0)}$ locally uniformly, we see that $d_{\ell(t)}(x,y)=d(g(t))(\Phi_{t}(x),\Phi_{t}(y))$ satisfies $d_{\ell(t)}\to(\Phi_{0})^{*}d_{g(0)}=:d_{0}$ locally uniformly. The completeness may be seen as followed: For fixed $t$ let $(x_{i})_{i\in\mathbb{N}}$ be a Cauchy sequence with respect to $d(\ell(t))$. We have, $d_{\ell(t)}(x_{i},x_{j})\leq\epsilon(N)$ for all $i,j\geq N$ where $\epsilon(N)\to 0$ as $N\to\infty.$ Hence $d_{g(t)}(\Phi_{t}(x_{i}),\Phi_{t}(x_{j}))\leq\epsilon(N)$ for all $i,j\geq N$. Since $g(t)$ is $a$-fair to $h$, we know that $g(t)$ is complete. Hence, $\Phi_{t}(x_{i})$ converges with respect to $g(t)$ to an $p$ as $i\to\infty$ which we can write uniquely as $p=\Phi_{t}(q),$ since $\Phi_{t}$ is a smooth diffeomorphism. Hence $(x_{i})_{i\in\mathbb{N}}$ converges, as $i\to\infty$, with respect to $\ell(t)$ to $q$. ∎
@@
Let $(M,g(0))$ be $C^{0}$ and $h$ be fixed, and satisfy $\sup_{M}|\nabla^{j}{\rm Rm}(h)|_{h}=:K_{j}<\infty.$ Let $\hat{\epsilon}\leq\epsilon(n)$ satisfy $\frac{1}{(1+\hat{\epsilon})}h\leq g(0)\leq(1+\hat{\epsilon})h$ and let $g(t)_{t\in(0,T)}$ be the solution to (3.1) from Theorem 3.1. Then there exists $\hat{T}(n,h,\hat{\epsilon})>0,\hat{T}\leq T$ such that
$|{\rm Rm}(g(t))|\leq\frac{\psi(n,\hat{\epsilon})}{t}$
for all $t\in(0,\hat{T}(n,h,\hat{\epsilon})),$ where $\psi(n,\hat{\epsilon})\to 0$ as $\hat{\epsilon}\to 0$. Hence the same is true for any Ricci flow related solution. In particular, if $inj(M,h)>0,$ then the Ricci flow related solution obtained in Theorem 3.4 satisfies this estimate also:
$|{\rm Rm}(\ell(t))|\leq\frac{\psi(n,\hat{\epsilon})}{t}$
for all $t\in(0,\hat{T}(n,h,\hat{\epsilon})),$ where $\psi(n,\hat{\epsilon})\to 0$ as $\hat{\epsilon}\to 0$.